A multiple choice examination contains 5 questions, each of which 4 options/choices (a,b,c,d). Suppose a student answers by guessing only, what is the probability that
a. he will get perfect?
b. his answer to all questions are of the same letters?

Answer by stanbon(75887)

A multiple choice examination contains 5 questions, each of which 4 options/choices (a,b,c,d).
Binomial Problem with n = 5 and p(correct) = 1/4
-------------------
Suppose a student answers by guessing only, what is the probability that
a. he will get perfect?
P(x = 5) = (1/4)^5 = 0.0009766
-------------------------
b. his answer to all questions are of the same letters?
Ans: (1/4)^5 = 0.0009766
==================
Cheers,
Stan H.